### v2

# Visual Web Agents with Realistic Tasks

[Uncaptioned image] WebGym: Scaling Training Environments for
Visual Web Agents with Realistic Tasks
Hao Bai1,2Alexey Taymanov1
Tong Zhang2Aviral Kumar3Spencer Whitehead1
1 [Uncaptioned image] Microsoft  2 [Uncaptioned image] UIUC  3 [Uncaptioned image] CMU
Work partly done during internship at Microsoft.
(January 7, 2026)
[…]
We present WebGym, the largest-to-date open-source environment for training realistic visual web agents. Real websites are non-stationary and diverse, making artificial or small-scale task sets insufficient for robust policy learning. WebGym contains nearly 300,000 tasks with rubric-based evaluations across diverse, real-world websites and difficulty levels. We train agents with a simple reinforcement learning (RL) recipe, which trains on the agent’s own interaction traces (rollouts), using task rewards as feedback to guide learning. To enable scaling RL, we speed up sampling of trajectories in WebGym by developing a high-throughput asynchronous rollout system, designed specifically for web agents. Our system achieves a 4-5x rollout speedup compared to naïve implementations. Second, we scale the task set breadth, depth, and size, which results in continued performance improvement. Fine-tuning a strong base vision-language model, Qwen-3-VL-8B-Instruct, on WebGym results in an improvement in success rate on an out-of-distribution test set from 26.2% to 42.9%, significantly outperforming agents based on proprietary models such as GPT-4o and GPT-5-Thinking that achieve 27.1% and 29.8%, respectively. This improvement is substantial because our test set consists only of tasks on websites never seen during training, unlike many other prior works on training visual web agents.
1 Introduction
Vision-language models (VLMs) have recently demonstrated remarkable performance on agentic tasks on the web. These tasks often resemble those that a human user would aim to perform on the web, such as “Compare the prices and specs between AirPod 3 and AirPod 2 on Apple.com”. To solve such tasks, an agent will perform a series of actions, such as navigating to the relevant website, searching for both products, and finally, deciding on and submitting an answer to complete the task. Some web agents rely purely on textual information and try to solve tasks by simply utilizing accessibility trees Zhou et al. (2023) as the observation for the agent, while we focus on visual web agents, where agents observe screenshots, much like humans. This choice is based on the premise that web interaction semantics are defined by the rendered interface users see, and that reasoning over visual affordances enables more robust generalization across diverse and dynamically generated web pages (Koh et al., 2024; Zhou et al., 2024; Shen et al., 2025).
There is a growing body of work on evaluating visual web agents by using benchmarks that involve either artificial or real-world websites. Benchmarks with artificial websites typically provide source code for hosting simulated environments that resemble real websites, such as REAL Bench (Garg et al., 2025) and VisualWebArena (Koh et al., 2024). In contrast, realistic benchmarks such as GAIA-Web (Mialon et al., 2023), BrowseComp (Wei et al., 2025), and InSTA-v3 (Trabucco et al., 2025) use tasks based on live websites. These “live-website” benchmarks tend to be more challenging, as real websites are inherently non-stationary (Bai et al., 2024, 2025a). For example, each time an agent visits the homepage of an e-commerce site, the set of displayed products may differ, and the same sequence of past actions could yield very different outcomes.
Recent work has demonstrated that vision-language models can be adapted into web agents via supervised fine-tuning and task-specific training pipelines (Hong et al., 2025). However, in contrast to the rapid progress of online reinforcement learning (RL) in text-only domains such as software engineering and mathematical reasoning (Pan et al., 2024a; Setlur et al., 2025), scaling RL for visual web agents remains substantially more challenging. Unlike text-based settings, where rollouts are fast and rewards are easily verifiable, visual web tasks require reasoning over fine-grained rendered interfaces and often lack unambiguous success signals (Huang et al., 2025a; Sarch et al., 2025; Xue et al., 2025), making data collection slow and inefficient. These challenges have inhibited scaling up post-training methods for visual web agents, particularly for diverse and long-horizon tasks, beyond current short-horizon settings (Zhou et al., 2024).
Thus, to train advanced visual web agents, we first need a strong task set that includes problems which are challenging for existing models, and that span diverse domains and goals to encourage generalizable policy learning. A good task set should be able to support scaling performance by scaling these dimensions. These tasks must also be paired with clear evaluation protocols that provide meaningful learning signals. Such a task set would need to be substantially larger and more varied than existing ones (Zhou et al., 2024). In addition, utilizing a large task set requires a high training speed, with rollout efficiency being the main bottleneck. Rollouts in visual web environments typically take far longer than the learning updates themselves, largely due to the high computational overhead of simulating web browsers in current RL pipelines.
[…]
2 Related Work
Visual web agents for realistic tasks. This work focuses on visual web agents designed to solve real-world tasks using vision-language models (VLMs). Such agents emulate how humans interact with browsers on websites: they take website screenshots and metadata as input, and generate interactive actions supported by the browser (e.g., clicking or typing) to make progress toward a given goal. This process repeats until the agent outputs an action indicating task completion, resulting in multi-step trajectories. Prior to this work, both general-purpose VLMs and web-specific visual agents have shown promising capabilities on this problem of training visual web agents. General-purpose VLMs can perform reasonable grounding and reasoning on web tasks when combined with prompt engineering or light post-training on web-related data to improve general agent behavior, as demonstrated by models such as Gemma3 (Gemma-Team et al., 2025), GLM-4.1v (Hong et al., 2025), and Qwen3-VL (Bai et al., 2025b). In contrast, web-specific visual agents are models that are heavily post-trained on web/web-adjacent tasks, like UI-TARS (Qin et al., 2025).
Evaluation benchmarks and training environments for realistic web agents. Recent work has made substantial progress in developing robust evaluation benchmarks for realistic web-browsing agents, while training environments remain relatively scarce. Existing benchmarks typically feature small, curated task sets with evaluation hints to assess specific reasoning or interaction abilities. Representative visual web benchmarks include WebVoyager (He et al., 2024), Mind2Web-Live (Pan et al., 2024b), and Mind2Web-2 (Gou et al., 2025), alongside broader but non-visual ones such as AgentSynth (Xie et al., 2025), InSTA (Trabucco et al., 2025), BrowseComp (Wei et al., 2025), TravelPlanner (Xie et al., 2024), DeepShop (Lyu et al., 2025), and GAIA (Mialon et al., 2023). A training environment, in contrast, must provide not only large-scale task sets but also efficient rollout and training systems. Efforts like WebRL (Qi et al., 2024) and the Proposer-Agent-Evaluator (PAE) framework (Zhou et al., 2024) have taken steps in this direction, but their task sets remain limited in diversity and hardness [^1] In this paper, “difficulty” or “difficulty level” specifically means the difficulty levels we propose as an attribute of the task set we propose, while we use the word “hardness” for indicating whether a given task is difficult or not., and their rollout systems are not optimized for speed. WebGym addresses these issues by expanding task diversity via rubric-derived fact groups and controlled decomposition into atomic and compositional variants (while limiting additional synthesis for sources like PAE-WebVoyager and AgentSynth to avoid near-duplicate synthetic goals), and by using a fully asynchronous rollout system that accelerates data collection and enables scalable reinforcement learning for web agents.
Scaling online reinforcement learning for training agents. Prior work has improved the performance of LLMs and VLMs by scaling the size and diversity of supervised fine-tuning (SFT) datasets, such as Super Natural-Instruct (Wang et al., 2022) and OpenThoughts (Guha et al., 2025). More recently, studies have shown that increasing the number of training trajectories through online RL can further boost agent capabilities, even with algorithms like REINFORCE (Williams, 1992), across software engineering (Pan et al., 2024a), math reasoning (Setlur et al., 2025), deep research (Zheng et al., 2025; Jin et al., 2025; He et al., 2025), and general reasoning. WebGym builds on these insights by constructing a large, diverse task set spanning multiple domains and difficulty levels, and training agents via on-policy RL from collected rollouts.
3 Tasks in WebGym and Evaluation Protocol
Training with online RL requires a large number of model-generated rollouts guided by reliable reward signals on a family of tasks (“environments” in traditional multi-task RL nomenclature) that we curate. To learn generalizable policies with such a procedure, the task set must (1) span diverse domains and websites, (2) cover a range of “difficulties” from atomic to compositional tasks, and (3) include verifiable evaluators that translate outcomes into meaningful learning signals. Existing web agent benchmarks, while effective for evaluation, are typically too small and static for large-scale training. Existing training tasks are either non-diverse (Zhou et al., 2024) and lack a wide range of task difficulties (Trabucco et al., 2025), or do not come paired with good task-level evaluation criteria, rubrics, or guidance for enabling robust policy evaluation (Trabucco et al., 2025; Zhou et al., 2024; Xie et al., 2025; Wei et al., 2025; Mialon et al., 2023). We describe how we procedurally construct a task set that satisfies these desiderata, with assistance from LLMs.
We start from a seed collection of high-quality tasks drawn from existing benchmarks and training environments (§3.1). We systematically annotate these tasks with evaluation criteria (evaluation references) and then decompose them into subtasks (§3.2) to achieve both breadth (coverage across many websites and domains) and depth (multi-criterion compositional tasks) (§3.4). The construction is fully done with the GPT-4o model and all prompts are provided in §E to ensure reproducibility and enable scaling up further. We then detail how we split the task set into train and test subsets (§3.3) and present our evaluation protocol (§3.5).
Source Task Set Difficulties # Websites # Tasks
InSTA-v3 (Trabucco et al., 2025) 146,348 146,441
PAE-WebVoyager (Zhou et al., 2024) 13 128,499
AgentSynth-Web (Xie et al., 2025) 328 2,086
BrowseComp (Wei et al., 2025) 7 1,266
TravelPlanner (Xie et al., 2024) 1 1,225
Mind2Web-Live (Pan et al., 2024b) 76 542
Online Mind2Web (Xue et al., 2025) 139 300
DeepShop (Lyu et al., 2025) 1 150
Mind2Web-2 (Gou et al., 2025) 44 130
GAIA-Web (Mialon et al., 2023) 1 87
WebGym (Ours) 127,645 292,092
Table 1: Task sourcing for WebGym. We aggregate web agent task sets from 10 widely-used benchmarks and environments to seed our procedural construction. We report original statistics (#difficulties, #websites, #tasks) for each seed set used to construct the WebGym task set. Some task sets are marked with only 1 website because the original task sets do not provide task-specific website.
1:Let $C$ be the task set with NN tasks, each having task description, difficulty, domain, website, and evaluator rubric fields.
2:Gather all tasks from datasets included in WebGym.
3:for each task in $C$ do
4:  Infer website and domain from task description.
5:  Generate evaluator rubric as fact groups using LLM, where each group contains one or more facts.
6:  Set difficulty as total number of facts.
7:  if at least 22 groups exist and at least one group has 33 or more facts then
8:   for each proper subset of groups containing at least one large group do
9:     Create decomposed task from subset using LLM, set its rubric and difficulty, then add to $C$.
10:   end for
11:  end if
12:end for
Algorithm 1 Task Set Construction Procedure of WebGym.
3.1 Obtaining a Seed Set of Tasks
To obtain a set of seed tasks, we collect 10 of the most recent and highly utilized task sets in the open-source community, as listed in Table 1. These tasks come from both existing benchmarks and existing training sets. Importantly, we do not designate any entire benchmark as “held out”, because many benchmarks share websites, domains, or task patterns with one another, making train-test separation at the level of existing benchmarks unreliable. Instead, WebGym constructs its own train-test split at the task and website level to ensure strict separation without discarding large amounts of useful data. Specifically, our tasks come from a) InSTA-v3 (Trabucco et al., 2025), which provides wide coverage of websites and domains; b) PAE-WebVoyager (Zhou et al., 2024) and AgentSynth (Xie et al., 2025), which provide synthetic tasks focused on narrow domains; c) BrowseComp (Wei et al., 2025) and GAIA-Web (Mialon et al., 2023), which include difficult tasks with concrete, verifiable answers; d) TravelPlanner (Xie et al., 2024) and DeepShop (Lyu et al., 2025), which focus on domain-specific browsing problems; and e) Mind2Web-Live (Pan et al., 2024b), Online Mind2Web (Xue et al., 2025), and Mind2Web-2 (Gou et al., 2025), which contain human-verified high-quality tasks that are indeed solvable. We next describe how we construct our full task set.
Figure 2: Task decomposition system. WebGym decomposes tasks by generating valid combinations of fact groups from the original task’s rubric. Decomposition requires $\geq$2 groups with at least one “large” group ($\geq$3 facts). Each valid combination (excluding the full set) produces a new task with lower difficulty while maintaining consistency with the original objectives.
3.2 Task Set Construction Procedure
Recent research has demonstrated strong improvements when constructing curriculum on tasks with similar goals but different hardness (Shen et al., 2025; Setlur et al., 2025). In parallel, other efforts show that developing evaluation rubrics can significantly improve evaluation accuracy (Pathak et al., 2025; Huang et al., 2025b; Gunjal et al., 2025). We combine these ideas by generating evaluation rubrics and using them to construct decomposed tasks. We elaborate our task construction algorithm in Algorithm 1 and an example in Figure 2. All constructions are generated using GPT-4o.
We begin by sourcing and merging all tasks specified in §3.1. For benchmarks that do not specify a concrete website for a task, we prompt GPT-4o to infer an appropriate website. Next, we use GPT-4o to generate an evaluation rubric structured as fact groups, where each group contains one or more evaluation facts. We define the difficulty or difficulty level of a task as the total number of facts across all groups in its evaluation rubric. The prompt we use to generate the evaluation rubric is shown in §E.5.
With these fact groups, we can procedurally generate decomposed tasks through combinations. Task decomposition occurs only when two conditions are met: (1) the original task has at least 2 fact groups, and (2) at least one group is “large” (i.e., it contains 3 or more facts). When these conditions are satisfied, we generate new tasks by selecting proper subsets of the fact groups, where each subset must contain at least one “large” group to ensure meaningful task complexity. For example, as illustrated in Figure 2, the seed task “Examine all the items displayed on the webpage, compare their prices and amounts and add the item with lowest unit price and one to shopping cart” contains 2 fact groups, with one fact group containing 3 criterion while the other fact group containing only 1. In this case, only the first fact group will participate in the decomposition, and thus our algorithm generates only 1 new task with the only “large” group.
Take a harder task for example, such as: “Find an eligible upcoming piano concert in the US or Canada by a Chopin Competition prize winner, providing the pianist’s name, competition year/prize, YouTube link to their final performance, and the concert’s date, city, venue, and event page link.” We decompose this into three fact groups: G1 (concert eligibility) with 2 facts (timing and location constraints), G2 (pianist details) with 3 facts (name, competition prize, YouTube link), and G3 (concert details) with 4 facts (date, city, venue, event link), yielding a total difficulty of 9. Since the decomposition conditions are met (3 groups $\geq$ 2, and G2 and G3 are “large” with $\geq$3 facts each), we generate valid subset combinations: {G2} (d=3), {G3} (d=4), {G1,G2} (d=5), {G1,G3} (d=6), and {G2,G3} (d=7). The combination {G1} alone is invalid since G1 has only 2 facts (no large group), and {G1,G2,G3} is excluded as it equals the original task. Example rubrics can be found in §D.1. We show the prompt we use to generate the decomposed tasks in §E.6.
As a result, tasks generated by WebGym (1) avoid being trivial by restricting to sufficiently large groups, (2) are strictly easier than the original task, yielding denser reward signals, and (3) are guaranteed to be well-defined whenever the original task is well-defined. In contrast, PAE (Zhou et al., 2024) generates synthetic tasks that may be ill-posed or impossible to solve. While WebGym is related to the approach in AgentSynth (Xie et al., 2025), it does not require executing agent rollouts to generate new tasks making it cheaper.
3.3 Train-Test Split
Having expanded the source tasks into a substantially larger task set, we next describe the train-test split. As we’re primarily interested in imbuing agents with broadly generalizable skills and capabilities, we construct an out-of-distribution (OOD) task set with 1,167 tasks where each OOD task originates from a distinct website. We remove all tasks in the training set associated with the same websites as those in the test set. This guarantees that all OOD tasks come from entirely unseen websites relative to the training set.
3.4 Statistics: Size, Breadth and Depth
Figure 3: Statistics of the WebGym training task set. Left: website distribution as a function of the sorted index (from more tasks to less) of website. Right: difficulty distribution of the train and test task sets. The transparent bars over the original bar mean decomposed tasks, and the slashed bars means test set.]
Figure 4: Analysis of the WebGym task set. Left: distribution of tasks across domains according to the Mind2Web-2 taxonomy (Gou et al., 2025). Right: Distribution of trajectory lengths by task difficulty for answered trajectories over multiple iterations. For trajectories working on medium- and hard-difficulty tasks that exceeds 30 tasks during evaluation, we filter in only trajectories under 30 tasks to make the comparison fair. The violin plots show that the probability density of trajectory lengths, with the KDE mode (black line, KDE mode is the peak of the smoothed probability density curve, where the violin is widest) indicating the most likely length and mean (red) indicating the average.
After the train-test split, the WebGym training set contains 292,092 tasks spanning 127,645 websites. The top 20 websites account for about half of all tasks (Figure 3, left), most of which are from the PAE-WebVoyager synthetic task set, while there are a significant number of websites with few tasks that build the diversity up, which is majorly composed of other source task sets. This reflects an intentional balance: many long-tail websites ensure cross-domain diversity, while frequent websites provide sufficient repetition for effective RL training. The task difficulty distribution (Figure 3, right) shows that around 80% of the tasks are Easy tasks (difficulty 1-3), which is intentional to encourage learning generalizable basic skills while we focus on more specific objectives at higher difficulty levels. To verify topical coverage beyond high-frequency sites, we compute domain diversity using the Mind2Web-2 taxonomy (6 domains, 24 subdomains) with GPT-4o. Figure 4 (left) illustrates broad domain coverage, with rollouts from different domains to promote diverse training samples. The trajectory length analysis (Figure 4, right) reveals a clear correlation between task difficulty and the number of steps required: easy tasks (difficulty 1-3) require 7.8 steps, medium tasks (difficulty 4-6) require 9.9 steps, and hard tasks (difficulty 7+) demand 11.9 steps on average. The KDE mode (which measures most likely trajectory length) also increases with difficulty, from 3.7 to 4.9 to 5.5 steps. This trend confirms that the tasks classified to be ”harder” in WebGym do actually involve more complex multi-step action and navigation before the agent can provide an answer.
3.5 Evaluator and Reward Function Design
Figure 5: Agreement between automated evaluators and human judgment. Rubric-based evaluation (with explicit criteria) consistently improves agreement over task-only evaluation, yielding higher accuracy and precision across LLM-based evaluators. Among the evaluators, GPT-4o shows the largest shift after adding the rubric: precision increases the most while recall drops, indicating that the rubric makes GPT-4o apply stricter, more conservative pass criteria.
Evaluating long-horizon visual agents is challenging because the produced trajectories often make partial progress, end up at visually similar but incorrect pages, and generate answers that may appear correct unless verified against grounded evidence. Unlike math reasoning or coding, “reference” answers of any sort are rare to obtain for most task instructions because only a small minority of benchmarks provide them (for example BrowseComp and GAIA-Web (Wei et al., 2025; Mialon et al., 2023)) and the notion of a reference answer does not exist in many cases (e.g., tasks that do not involve question answering). This difference reflects a fundamental challenge in designing open-ended web tasks with a binary answer. Thus, evaluation cannot rely on direct matching to ground-truth outputs and must instead interpret the trajectory in context. To this end, we mainly focus on LLM-judged evaluation, while reference answers, when available in the source benchmark, overrides the judge.
For faithful evaluation, recent research has adopted a criterion proposal approach that generates task-specific criteria to guide evaluation (Kimi-Team et al., 2025; Chen et al., 2025). However, we posit that tasks usually come with structures, so we propose to generate rubric with structures such as fact groups. As discussed in Section 3.2, in WebGym, each task is paired with a rubric of several fact groups, where each group contains one or more criteria that guides evaluation reducing false positives from ambiguity about task success.
With the task-specific rubric, a trajectory receives reward only when all criteria are satisfied, producing a binary signal consistent across domains and difficulty levels. Since not all screenshots are informative, the judge focuses on evidence-bearing pages, obtained via a high-recall keypoint selection process inspired by Online Mind2Web (Xue et al., 2025) that retains relevant observations while filtering out distractors such as detours and ads. In practice, we use GPT-4o for both keypoint selection (prompt shown in §E.2 and per-criterion evaluation (prompt shown in  §E.3). The evaluator proposes keypoint screenshots from the task description, iterates through trajectory steps, and issues binary judgments for each criterion using the selected keypoints and final answer.
We validate the evaluator against human annotations on 80 collected trajectories sampled uniformly across difficulty levels (10 trajectories per level from difficulties 1-6, and around 5 trajectories per level from difficulties 7-10), where nearly half of the trajectories in each difficulty are marked as correct by the WebGym evaluator to balance the set, and find that rubric-guided evaluation improves agreement over task-only judging, which increases accuracy and precision for all evaluator models being tested (GPT-4o, Qwen3-VL-8B-Instruct, and Gemma3-27B-it). We observe minor regressions in the recall metric for the subset of stronger evaluator models (Qwen3-VL-8B-Instruct and especially GPT-4o), where some rubrics tend to be overly strict (Figure 5). Note that this regression is almost impossible to eliminate when using an LLM-based rubric proposer: depending on the prompt used to instruct the model, it may generate stricter rubrics, occasionally resulting in over-strict evaluations. Examples of evaluations with and without criteria are shown in §D.2. This trade-off is ultimately beneficial for RL training: we exchange a slight reduction in sample efficiency for more precise but conservative learning signals, but at the same time, this helps prevent error accumulation in long-horizon settings.
4 Rollout System in WebGym
The primary bottleneck in scaling up RL training for web agents lies in the slow speed of rollout generation, particularly in multi-step RL settings. This challenge becomes more severe when simulation resources, such as CPUs, are limited. To make WebGym practical for large-scale agent training, we address this issue by replacing synchronized, batch-style rollout collection with an asynchronous system that keeps both CPUs and GPUs fully utilized throughout multi-step rollout generation. We show in Figure 6 how tasks are queued into a process pool, how stateful browser sessions are isolated for each rollout, and how new tasks immediately begin inference once resources become available, eliminating the need for global synchronization barriers. We now describe this system and benchmark its performance and throughput.
To our knowledge, the WebGym provides the the first, open-source rollout system optimized for asynchronous multi-turn simulations for web agent tasks. In contrast to work on RL infrastructure for LLMs, WebGym focuses on an orthogonal dimension: speeding up rollout generation up by introducing a highly efficient simulation server for web browsing so that the rollout pipeline does not synchronize between any step or episode. This server can be easily integrated with any RL framework that focuses on system-level speedup like optimizing the weights save/load overhead and algorithmic speedup like introducing episode/step/token-level off-policy RL. Examples of these RL frameworks include veRL (Sheng et al., 2025), veRL-agent (Feng et al., 2025), PipelineRL (Piché et al., 2025), SkyRL (Cao et al., 2025), and AReaL (Fu et al., 2025).
We will use the simplest implementation of the RL loop in our experiments, where rollout and training happen synchronously. We instead focus on optimizing rollout generation as discussed above. Since we only optimize rollout generation, in this section, our benchmarking numbers are based on inference-only experiments, and no training is involved.
4.1 Strategy and System Architecture
Naïve implementations of rollout systems are synchronous in nature, like what was implemented in Shen et al. (2025); Bai et al. (2024); Qi et al. (2024), where a fixed group of browser sessions is forced to advance at the same pace to perform some number of rollouts in parallel. At each step, every session loads its webpage on the CPUs, the system sends all screenshots from the group together to the GPU for one combined policy call, and then waits until the slowest session completes before any session proceeds; a common variant waits at the end of each full trajectory instead of at every step. These step and episode barriers are problematic in web settings because step times and horizons vary widely across tasks and even across data-collection policies. Faster sessions sit idle while slower ones catch up, and hardware activity arrives in synchronized bursts that saturate either CPUs or GPUs at once, and then drop to idle. We illustrate this “burst-idle” behavior in Figure 6 and then quantify its impact on end-to-end data collection time and scaling behavior in Figure 7.
Figure 6: Asynchrony eliminates burst-idle behavior in web rollouts. Left: WebGym implements an asynchronous rollout system that (1) shortens single-trajectory collection duration by isolating rollout processes and (2) allows new rollout processes to join early by replacing batched inference with a process pool; the example shown is a toy case with 3 available environment buckets but 4 tasks waiting to roll out. Right: CPU utility trace over a sampled rollout period, comparing the synchronized rollout framework to WebGym: synchronized barriers induce spiky “all-busy then all-idle” utilization, whereas WebGym streams work as soon as observations are ready, smoothing CPU load and reducing idle gaps while CPU workers await policy actions.
Figure 7: Benchmarking speed and throughput of the WebGym asynchronous rollout framework. Left: the WebGym framework boosts the rollout speed up significantly with a 4x-5x speedup. This figure shows time cost and average CPU utility percentage when reaching all 256 environments running w.r.t. different amounts of CPUs, while using the same amount of GPU resources for running inference on the VLM-based agent. The limit of browsers per CPU can scale to 2 without crashing. Right: the WebGym framework boosts the rollout speed up linearly w.r.t. GPU nodes. This figure shows WebGym throughput w.r.t. different numbers of GPU nodes when enough CPUs are provided on each machine (thus speed is bounded by GPU).
To mitigate the inefficiency from this “burst-idle” behavior, our framework adopts a server/client architecture (Ali et al., 2020), where the server hosts simulation environments and the client organizes the rollout loop. The server simulates the browser environments following a master/worker paradigm (Aida et al., 2003), with a master node routing API requests to CPU worker nodes that host the actual simulators and execute actions. The master node configuration determines how much parallel work the system can support, making it easy to scale horizontally by adding workers or vertically by increasing per-node resources. The server is entirely CPU-based and requires no GPUs. The client hosts agent instances on the GPU and send requests to the server. To manage the requests sent to the server, we adapt an operation-specific local queue that enables spreading CPU and GPU workload evenly under tight resource settings, with more details in §B.1. The client is GPU-heavy and relies little on CPU budgets.
4.2 Benchmarking our Rollout System
We demonstrate the strong capability of the framework under different GPU/CPU resource constraints, comparing WebGym to a synchronized multi-step rollout strategy, as shown in Figure 7 (left). We fix the number of GPUs (to 3 nodes of 8 NVIDIA H100 GPUs) and environments for all experiments to collect 1800 trajectories with an average (instead of max) of $13.2$ steps thus in total 23,760 steps, then modify only the number of CPUs on the asynchronous rollout server. For all runs, we use CPU clusters where each worker node provides 64 AMD EPYC 7763 CPUs.
[…]
When CPU resources are scarce, the asynchronous design streams observations to the GPUs as soon as they are available, preventing GPU starvation and maintaining high GPU utilization while smoothing CPU load. This contrast between CPU-limited and GPU-limited regimes explains the utilization patterns and measured time-to-trajectory observed in Figure 7. We also benchmark the scalability of WebGym by increasing the number of concurrent trajectory collections. In this experiment, we provision sufficient CPU resources (768 CPUs) and vary the number of H100 GPU nodes used by the system. Figure 7 (right) shows the system exhibits strong scaling that is sub-linear, but very close to the desired linear speed-up curve. This slowdown primarily stems from the synchronization overhead from coordinating rollouts across many nodes.
5 Training Web Agents with WebGym
We train web agents using the tasks and system provided by WebGym. Our goal is to study the design choices behind a simple, visual web agent by training on WebGym. Concretely, we center our empirical results around two central questions. First, we study how WebGym can transform existing VLMs to efficiently solve complex web tasks via RL training (§5.2). Second, we investigate how the agent’s generalization capability scales when trained with the large and diverse task set provided by WebGym, which eventually leads to performance improvement (§5.3). We show that via a simple REINFORCE-style RL training procedure, we can train agents that improve performance from the initial 26.2% to 42.9% on the test split of WebGym. The hyper-parameters of the best run are shown in §F.
5.1 Experimental Setup
To study the two questions outlined above, we first establish the experimental setup we operate in.
Action space. We parameterize the action space by extending the definition of Shen et al. (2025) with a navigation action, so that the action space now includes clicking, typing, scrolling, returning to the previous page, and navigating to a website specified by the agent. The WebGym rollout system supports both the Set-of-Marks mode (Yang et al., 2023), where interactive elements are labeled in the screenshots by the simulator, and a coordinate-based screenshot-only mode. We operate directly in the coordinate-based mode, which is supported by the base model (Qwen3-VL-8B) we use to train via RL.
Policy update. We run REINFORCE (Williams, 1992) with binary terminal rewards, without any baseline or negative gradient. This update is equivalent to online filtered behavior cloning (BC) (Bai et al., 2024) or a “thresholded” version of reward-weighted regression (Peters and Schaal, 2007), which retains only successful trajectories and maximizes their log-likelihood. The training objective is given by:
$\operatorname*{arg\,max}_{\theta}~~\mathbb{E}_{\mathcal{T}\sim\text{tasks}}\left[\mathbb{E}_{o_{0:\tau},\,a_{0:\tau-1}\sim\pi_{\theta}(\cdot\mid\mathcal{T})}\left[\left(\sum_{t=0}^{\tau-1}\log\pi_{\theta}(a_{t}\mid o_{\leq t},\mathcal{T})\right)\cdot\mathds{1}\!\left[R(o_{0:\tau},\mathcal{T})=1\right]\right]\right],$ (1)
where $\theta$ denotes the parameters of the (multi-step) agent policy $\pi_{\theta}$, and $\mathcal{T}\sim\text{tasks}$ is a task instance sampled from the training task distribution. A rollout proceeds in discrete interaction steps indexed by $t$. At each step $t$, the agent samples an action $a_{t}\sim\pi_{\theta}(\cdot\mid o_{\leq t},\mathcal{T})$ conditioned on the full observation history $o_{\leq t}$ (and the task), and the environment returns the next observation. The rollout terminates at step $\tau$ only when one of the following two conditions are met: (i) when the agent emits an ANSWER action, or (ii) when the step index reaches the training horizon $h_{\text{train}}$, the maximum allowed number of interaction steps during training; thus $\tau\leq h_{\text{train}}$. The realized trajectory therefore contains actions $a_{0:\tau-1}$ and observations $o_{0:\tau}$, where $o_{0}$ is the initial observation and $o_{\tau}$ is the terminal observation after the last action. If termination occurs under the first condition (i.e., when an ANSWER action is produced), the environment does not advance and we set $o_{\tau}=o_{\tau-1}$ (i.e., the final observation is identical to the last pre-answer observation). The binary terminal reward $R(o_{0:\tau},\mathcal{T})\in\{0,1\}$ indicates whether the overall trajectory succeeds on task $\mathcal{T}$. The objective increases the log-likelihood of all actions along successful trajectories while assigning zero weight to unsuccessful ones, avoiding negative gradients from failures. In principle, we expect that incorporating a baseline (negative gradient) would only improve performance further, but chose to avoid it for now for faster experimentation due to improved training stability with only a positive gradient.
Base policy. We use the Qwen3-VL model family for its strong grounding and reasoning capabilities (Bai et al., 2025b; QwenLM-Team, 2025). We experiment on the 8B-sized model with both Instruct and Thinking variants.
Task sampling approach. We ablate on random task sampling and ratio-over-difficulty sampling over the WebGym task set, and present results for both approaches in §5.3. We also maintain a dynamic blocklist of websites that prevent access, avoiding those sites for subsequent training periods (§B.2).
Controlling horizon during training. To keep rollout collection efficient, we cap the maximum number of interaction steps per episode. Although these trajectories may still be successful (i.e., they eventually solve the task), they can be suboptimal in terms of step count, taking many unnecessary interactions compared to a shorter solution. Without a step cap, rollouts may include many such long-but-successful trajectories, which are expensive to collect and dilute training with avoidable steps. By truncating the horizon, we reduce wasted environment interactions during data collection and bias training toward shorter, more efficient successes, improving step efficiency. This is also preferable to simply tuning the learning objective to emphasize trajectories shorter than a threshold while still collecting long trajectories, because the horizon cap prevents spending those extra steps in the first place rather than only downweighting them after collection. We study the effect of the training horizon in §5.3.
Probing task hardness with proprietary VLMs. To understand the hardness of the OOD test set in WebGym, we also evaluate GPT-4o and GPT-5 (Thinking) as two strong proprietary models on the OOD test set. Results show that simply using the same prompt as Qwen3-VL (§E.1) with Set-of-Marks (SoM (Yang et al., 2023), used here because these models cannot accurately output coordinates) leads to non-trivial but far below 100% performance, where GPT-4o achieves 27.1% on the full test set, and GPT-5 achieves 29.8% on a randomly sampled subset with 300 tasks from the OOD test set (subsampled due to the budget costs associated with GPT-5). Note that GPT-4o achieves 25.6% on this task subset, so the variance caused by subsampling is not significant. As we discuss next, our trained agent significantly outperforms both of these compared approaches.
Refer to caption(L) Refer to caption(R)
Figure 8: Ablations on base models, prompting, and action filtering. (left) test-set success rate curves of different models (Qwen3-VL-Instruct-8B, Qwen3-VL-Thinking-8B, GPT-4o, and GPT-5-Thinking), and of Qwen3-VL-Instruct-8B under different constraints (either removing the memory prompt or removing the repetition penalty during RL). (right) test-set same-screenshot (repetitive inefficient action) rate and trajectory length curves before and after applying the action repetition penalty (filtering) on the Instruct model. The orange curve is an average of two runs, while the translucent orange area around the orange line denotes the variance of these two runs. We observe that even if the starting variance between different runs can be large, the curves converges after 24k training trajectories are collected. All plots are produced with an EMA (Exponential Moving Average) smoothing of $0.50$.
5.2 Design Choices for Agent Training
In this subsection, we describe the key design choices we had to utilize to train agents via RL on WebGym. Inspired by prior research on RL for LLMs (Setlur et al., 2025; Xiong et al., 2025), we note that biasing the sampled tasks towards hard tasks improves the performance and sample efficiency, as long as easy tasks are also used either via co-training or via a prior stage of a curriculum. Thus, we start from a difficulty sampling that slightly biases towards harder tasks (concretely, easy : medium : hard = $2:5:3$). All experiments in this section follow this task sampling strategy, while we study different sampling strategies further in §5.3.
Memory prompt for long-horizon tasks. Our base model (Qwen3-VL-8B) follows a simple prompting template for computer-use agents, which includes a high-level action description (e.g., “click on the search button”) paired with a low-level tool call (e.g., “Click [209, 441]”) (Bai et al., 2025b; QwenLM-Team, 2025). The exact prompt is shown in §E.1. From preliminary experiments, we observe that this template yields strong initial performance, because the Qwen-3 series of models are trained on data collected from web environments with this prompt (Bai et al., 2025b; QwenLM-Team, 2025), but this prompt provides limited improvement if we use it to rollout trajectories with RL on WebGym, as shown in Figure 8 (left).
This limitation arises from the model’s inability to retain and reuse information gathered in earlier steps. For example, in the task shown in Figure 2, where the goal is to compare two products, the agent must recall details about the first product when evaluating the second. In the current design, this requires including the entire conversation history and all prior screenshots as input, which is inefficient and impractical. A more formal treatment of this challenge from a partially observable Markov decision process (POMDP) perspective is provided in §C. To solve this limitation, we propose that the model should maintain a memory by responding with the updated memory in each step, inspired by GLM-4.1v (Hong et al., 2025). The prompt with memory is also demonstrated in §E.1. Utilizing the memory prompt significantly boosts performance after RL (see Figure 8, left, Instruct-8B versus w/o Memory Prompt).
Removing repeated actions with a reward penalty. We frequently observe cases where the base model is stuck at the same screenshot and attempts to repeat the same action multiple times. In order to measure how often this behavior happens, we calculate the fraction of times the screenshot at one step is the same as the next one (% Same Screenshot), and the trajectory length (Avg. Steps per Trajectory) in Figure 8 (middle). Preliminary experiments show that standard RL substantially reduces this behavior of repeating ineffective actions, but does not fully eliminate it(Figure 8, middle). Thus, we design a penalty for these repeated actions when doing RL: we explicitly filter out a step if the resulting next step presents an identical screenshot, even if the current step is a part of a successful trajectory. After applying this penalty, our RL runs become substantially more sample efficient as shown in Figure 8, left (Instruct 8B vs “w/o Repetition Penalty”).
Use of thinking for every step. In addition to the Instruct variant of Qwen3-VL-8B, we also experiment with the Thinking variant of this model with exactly the same prompt as the Instruct model. Results as shown in Figure 8 (left) reveal that the Thinking model attains a higher initial performance compared to the Instruct model, suggesting that extended reasoning at each step can be beneficial. However, the Thinking model produces significantly longer responses: before RL training: for instance, it produces 2139 characters per response compared to 1088 of the Instruct model on the held-out test tasks. Thus training the Thinking model incurs a larger latency cost and requires more compute. Notably, as training progresses with WebGym, the performance gap between the Instruct and Thinking models decreases considerably, and soon the Instruct model surpasses the Thinking model. Given the favorable trade-off between performance and efficiency, we adopt the Instruct model for subsequent experiments, though we believe that developing a compute-efficient RL recipe to effectively train a Thinking model is an interesting direction for future work.
5.3 Main Results: Scaling Agent Performance with WebGym
Base Model Prompt RL on WebGym Performance (%)
Proprietary Models GPT-4o Memory $27.1$
GPT-5 Memory $29.8$
Open-source Models Qwen3-VL-Instruct-8B Official* $27.4$
Qwen3-VL-Instruct-8B Memory $26.2$
Qwen3-VL-Thinking-8B Memory $28.2$
Qwen3-VL-Instruct-8B Memory ✓(Ours) $\mathbf{42.9}$
Table 2: WebGym empowers the small Qwen3-VL-Instruct-8B model to learn with RL to achieve state-of-the-art performance on the holdout test task set, arriving at a strikingly 42.9% success rate, significantly surpassing all agents empowered by proprietary models and small-yet-effective open-source agentic models. *The Official prompt refers to the web agent prompt specified in QwenLM-Team (2025).
Next, we present our main results scaling agent performance with WebGym. In particular, we are interested in understanding what concretely about the design of WebGym scales performance the most, and how much improvement we can obtain by running RL on WebGym. In the end, we’re able to scale the performance up to 42.9%, significantly surpassing existing agents empowered by proprietary models like GPT-4o and GPT-5-Thinking and strong open-source models like Qwen3-VL-Instruct-8B and Qwen3-VL-Thinking-8B. The comparison of these results are demonstrated in Table 2. In this section, we will understand how WebGym enables these results by specifically studying the effects of: (i) breadth (domain diversity), (ii) depth (difficulty composition), and (iii) the train-time interaction horizon, during RL training on WebGym. Our starting point is the best checkpoint from §5.2, which biases towards harder tasks (easy : medium : hard = $2:5:3$).
1) Scaling task set breadth (domain coverage) for better generalization. To study the effect of having a wide variety of domains in WebGym, we randomly remove half of the subdomains from the training set (“exclude domains”) while keeping the same training recipe otherwise as all the comparisons. This results in the subsampled training task set that only contains 53% of original tasks. Across all evaluation slices, removing domains (“exclude domains”) compared to using all tasks from WebGym (“biased to hard”) consistently slows down improvements and lowers the final success rate (Figure 9a–d), indicating that domain breadth matters even when the difficulty mixing ratio is unchanged. This gap is not due to oversampling and overfitting on fewer tasks: note that the number of collected training trajectories (36k) is far smaller than the number of training tasks remaining ($\sim$ 150k), so most training happens on distinct tasks. The degradation reflects reduced domain diversity, which limits generalization to unseen websites and subdomains.
Figure 9: Exploring scaling dimensions of WebGym and difficulty-aware training. Test-set success rate curves under different variations to training: removing domains (“exclude domains”), tuning difficulty ratios (“uniform sampling”, “biased to hard”, “only easy”, “only medium”), and shortening the train-time step budget (“shorten horizon”). Results are reported for (a) Overall (all tasks), (b) Easy (1–3), (c) Medium (4–6), and (d) Hard (7+). The orange curve averages two runs; the translucent orange band shows their variation. All plots are produced with an EMA (Exponential Moving Average) smoothing of $0.50$.
To study 2) and 3) below, we compare four task compositions: a) training only on easy tasks (“only easy”, $1:0:0$), b) only on medium tasks (“only medium”, $0:1:0$), c) a task mixture biased towards harder tasks (“biased to hard”, $2:5:3$, as used in §5.2), and d) sampling tasks uniformly at random from the full training set (“uniform sampling”, approximating a ratio of $25:5:1$ given the task set).
2) Scaling candidate task set size by sampling more easy tasks avoids overfitting to narrow hard tasks with the filtered BC loss. In our previous experiments, we sample with a difficulty mixture of $2:5:3$, which we note is heavily biased towards harder tasks compared to uniform sampling, because easy tasks actually comprise around $80\%$ of all tasks. This can make the effective task set size (i.e., expected number of tasks sampled at any given training iteration that are distinct from all current and previous iterations) very small because a large portion of the training tasks are repeated high-difficulty tasks. As we roll out more and more trajectories, we believe that this imbalance leads to overfitting because the test performance curve begins to drop right around 36k-42k training trajectories on the x-axis, as shown in Figure 9 (biased to hard). This problem can, in principle, be resolved with a better loss function or, perhaps more simply, be resolved by scaling the size of candidate task set. We choose to scale up the size of the candidate task set by adjusting the task sampling ratio so that the sampled tasks include more easy tasks that are distinct from each other.
In fact, we found that training exclusively on easy tasks performs surprisingly well and is more stable than over-emphasizing medium and hard tasks. In particular, the “only easy” setting consistently outperforms “only medium” on overall success as well as on easy and medium test tasks (Figure 9a–c), and it does not exhibit the late-stage plateauing or mild performance regression observed when harder tasks are upweighted (e.g., the “biased to hard” setting after $\sim$36k trajectories; Figure 9a–c). This behavior contrasts with typical findings in LLM reasoning, but is consistent with the structure of WebGym, where the very large set of easy tasks spans many websites and domains and thus supports generalization to unseen websites. On the other hand, very hard tasks in math reasoning can be built for all sub-areas for mathematics. While repeated exposure to a narrower set of domains that provide harder tasks remains important (as discussed below), the size afforded by easy tasks plays a crucial role in mitigating overfitting and enabling generalization to new websites. In contrast, aggressively upweighting medium and hard tasks increases repeated exposure to a smaller subset of domains and interaction patterns, leading to overfitting that manifests as performance plateaus and slight degradation in our results.
3) Scaling candidate task set depth (wider range of difficulty levels) by including tasks from all difficulty levels to promote more efficient learning. Despite the strong performance of the “only easy” setting, we believe that higher-difficulty tasks still play an important role. This intuition is also supported by our empirical observations: the performance gap between “only easy” and “biased to hard” is substantially smaller on the medium- and hard-difficulty test subsets than on the easy-difficulty subset. Motivated by this, we increase the depth of the task set by incorporating tasks spanning a wider range of difficulty levels. The simplest way to achieve this is through uniform sampling from the full training set. We find that uniform sampling yields the highest overall performance among the four difficulty-mixing strategies we evaluate (Figure 9a), and it achieves a significantly higher success rate on the medium-difficulty test set compared to training on easy tasks alone (Figure 9c). Because the “only easy” setting already draws from a large and diverse pool of tasks, the additional gains from uniform sampling are best attributed to the inclusion of some medium- and hard-difficulty tasks. Injecting these harder tasks provides training signals that enable the agent to develop capabilities beyond easy interactions, particularly on medium-difficulty test tasks. Finally, since uniform sampling includes only a small fraction of hard tasks, we do not observe a significant difference between these two settings on the hard-difficulty test set.
4) Horizon: controlling train-time interaction budget. Additionally, we study one important train-time hyperparameter for algorithmic exploration: the train-time interaction budget. To control this budget and shape the distribution of training trajectories, we reduce the per-episode step limit from $(15,30,45)$ to $(10,20,30)$ (“shortened horizon”). Empirically, shortening the horizon improves both sample efficiency and final performance across all difficulty slices. In particular, the success rate increases from a peak of 38.2% at the previous best checkpoint (uniform sampling with a step budget of $(15,30,45)$) to a peak of 42.9%. We hypothesize that a tighter step budget acts as a regularizer on the policy’s interaction process. By limiting opportunities for late-stage recovery, it discourages trajectories that succeed only after extended sequences of low-yield exploration, and instead concentrates training signal on earlier, higher-impact decisions. This increases the density of informative state–action transitions per episode, improving credit assignment and reducing variance in policy updates. Notably, performance on hard tasks also improves substantially (Figure 9d), consistent with the model learning more robust and reusable interaction primitives (e.g., efficient navigation and decisive goal completion) that transfer beyond the capped training horizon and support longer-horizon problem instances at evaluation time.
Setting Legend in Figure 9 Peak Performance (%) $\Delta$ from Previous (%)
Baseline Exclude Domains 31.0
+Breadth Biased to Hard 34.5 +3.5
+Size Only Easy 36.9 +2.4
+Depth Uniform Sampling 38.2 +1.3
+Horizon Control Shorten Horizon 42.9 +4.8
Table 3: Scaling overview: peak performance of different scaling directions through RL settings on the test set. $\Delta$ from Previous shows the improvement over the immediately preceding setting.
All main results from this section are summarized in Table 4, and the overview of the scaling results we discussed is presented in Table 3. Our final agent is trained with the memory prompt and penalty for repetitive actions, with a task horizon of $(10,20,30)$ over (easy, medium, hard) tasks, and with uniform sampling across all task difficulties, surpassing the best proprietary model evaluated (GPT-5-Thinking) by 13.1%, as shown in Table 2 and Table 4. We provide raw results to promote fair comparison, and encourage readers to include raw results in future work to compare with these scores.
6 Discussion and Conclusion
We introduced WebGym, the largest open-source training environment for visual web agents. Our procedural task construction approach generates structured evaluation rubrics and decomposes seed tasks into valid criterion subsets, enabling both domain breadth and hardness depth while ensuring decomposed tasks remain semantically coherent. To support large-scale training, we developed an asynchronous rollout system that eliminates synchronization barriers, substantially accelerating data collection for on-policy RL.
Our experiments reveal several empirical insights: explicit memory mechanisms are essential for long-horizon tasks requiring cross-step information retention; penalizing repeated ineffective actions improves sample efficiency; domain diversity directly translates to out-of-distribution generalization; uniform difficulty sampling outperforms both easy-only and hard-biased curricula; and shorter training horizons improve both efficiency and final performance. With these design choices, a simple REINFORCE algorithm enables strong generalization to entirely unseen websites, outperforming proprietary models.
Future work. Despite the availability of meaningful task-specific rubrics in our task set, LLM-generated rubrics can sometimes be overly strict, which slightly reduces sample efficiency during training. An important direction for future work is to explore more advanced designs for rubric-based evaluation. One promising approach is to treat LLM-generated rubrics as soft guidance rather than strict criteria, for example by training an evaluator that jointly conditions on both the rubric and the task instruction. Another promising direction is to investigate reinforcement learning algorithms that utilize WebGym more efficiently, particularly methods based on dynamic curricula (Shen et al., 2025), automatic curriculum learning (Bai et al., 2024), multi-agent training, and cross-site navigation.
Acknowledgements
We thank Rui Yang, Rui Pan, Yifei He from UIUC, Yifei Zhou from UC Berkeley, Anikait Singh, Salman Abdullah from Stanford for their early discussions of this research, members of the CMU AIRe lab for their feedback on human evaluation and presentations, and Eduardo Salinas, Gustavo De Rosa, Ahmed Awadallah from Microsoft Research (AI Frontiers Lab) for their continued support throughout this work.
Appendix A Raw Experimental Results
We demonstrate the raw values of all experiments presented in §5 in Table 4 below.
Method $H_{\text{train}}$ Sampling Domain Memory Rep. Penalty Zero-shot (%) RL Peak (%)
[…]
Thinking-8B $15:30:45$ $2:5:3$ Complete $28.2$
W/o Memory Prompt $15:30:45$ $2:5:3$ Complete $27.4$ $31.5$
W/o Repetition Penalty $15:30:45$ $2:5:3$ Complete $26.2$ $33.0$
Uniform Sampling $15:30:45$ $25:5:1$ Complete $26.2$ $38.2$
Only Easy $15:30:45$ $1:0:0$ Complete $26.2$ $36.9$
Only Medium $15:30:45$ $0:1:0$ Complete $26.2$ $35.0$
Exclude Domains $15:30:45$ $2:5:3$ Half $26.2$ $31.0$
Shorten Horizon $10:20:30$ $25:5:1$ Complete $26.2$ $\mathbf{42.9}$
GPT-4o-SoM $27.1$
GPT-5-SoM (Think) $29.8$
Table 4: Raw ablations and performance results of our trained agents on the OOD test set. $H_{\text{train}}$: training horizon (easy:medium:hard steps); Sampling: difficulty ratio (easy:medium:hard); Domain: subdomain coverage; Memory: memory prompt; Rep. Penalty: repetition penalty. For Instruct-8B (Biased to Hard), we report the peak value of the averaged curve from two runs, with $\pm$ indicating the half-range.
Appendix B Training Environment Designs
B.1 Rollout System: Operation-specific Local Request Queue
To avoid flooding the server with requests, the most naive approach is to implement a global queue (first-in-first-out rule) to avoid flooding the server with requests. However, below we show the ineffectiveness of this implementation because (1) it causes extensive warmup time and (2) it creates operation bottlenecks. To improve it, we demonstrate that each operation type should have its own queue. This is to say, requests with different types of operations should not have any priority over each other.
Figure 10: WebGym implements an operation-specific queue system that balances the CPU and GPU machine usages. Here we illustrate the computational model of this pool system in the left subplot, and the comparison of the two queue designs on the right. We represent the centralized CPU queue with Queue, and the cascading queue system for with the initial letter of the request type, specifically, Navigating to a webpage (N), taking a Screenshot (S), and Executing Action (E) that takes medium amount of time. A green box means the CPU machine is under optimal load.
The computational model of rolling out a trajectory in a multi-step RL setting is shown in Figure 10 (Left). When simulation is required for an RL environment, there will be a sequence of CPU operations bounded by CPU resources. On the other hand, high-speed RL inference usually put high pressure on GPU resources as well. For the web agents rollout collection, the computational model can be simplified to five stages: navigating to a website with the simulator (N), taking a screenshot with the simulator (S), proposing an action with the agent, executing an action with the simulator, and sending the trajectory to an evaluator to obtain reward (E), with their speed defined on the right side of each box.
Assume we send a burst batch of 180 requests to the rollout system at $t=0s$. As shown in Figure 10 (middle), with a global queue, at $t=1s$, although 60 Navigation requests are finished, the Screenshot requests arrive strictly after the rest of the Navigation requests, so they have to wait for all Navigation requests to return. This makes GPU starve at $t=2s$. Similarly, after $t=5s$, only after all Screenshot requests are returned can the Execution requests be processed, and after all Execution requests can be processed can the next-step Screenshot requests be processed, causing the GPU to starve again and again during the multi-step loop.
To improve this, WebGym implements a local queue that does not specify any priority among operations. At $t=1s$, Screenshot requests and Navigation requests will be processed with the same priority. This results in 30 Navigation requests and 15 Screenshot requests being returned at $t=2s$ (as they share CPU resources). When propagating this pattern through the pipeline, the GPU resources will not starve at any moment.
Figure 11: Left: Number of trajectories successfully finished and crashed out under high CPU load when $t=10\text{min}$ with 64 CPUs, 256 environments, and 3 GPU nodes. Top line: global queue. Bottom line: local queue. Right: The block rate increases as training goes on.
We experiment with the performance boost of this local queue under extreme CPU condition (64 CPUs needs to run 256 environments). Results are shown in Figure 11 (left). We observe that the crash percent decreases significantly if we employ the local queue, showcasing the effectiveness of this design.
B.2 Rollout Sampler: Anti-blocking
As we observe that some websites block frequent requests from the same server as shown in Figure 11 (right), we maintain a list of websites that blocks the agent with an LLM. The blocking is detected by GPT-4o with the prompt specified in §E.4.
Appendix C Modeling the Markov Decision Process On Long-Horizon Web Navigation Tasks
A partial observable Markov decision process (POMDP) extends a Markov decision process (MDP) to settings where the agent does not have direct access to the underlying state but instead receives incomplete observations. Formally, a POMDP is defined as a tuple $(S,A,P,R,O,\Omega,\gamma)$, where $S$ is the set of latent states, $A$ is the action space, $P(s^{\prime}\mid s,a)$ defines the state transition dynamics, $R(s,a)$ is the reward function, $O$ is the observation space, $\Omega(o\mid s)$ specifies the observation model that maps hidden states to observations, and $\gamma\in[0,1)$ is the discount factor. Because the agent does not directly observe $s_{t}$, it maintains a belief state $b_{t}$, a probability distribution over $S$, which it updates using past actions and observations. As a result, the optimal policy maps belief states to actions, $\pi(b_{t})\rightarrow a_{t}$, instead of mapping fully observed states to actions as in standard MDPs.
Recent work models the web agent tasks as a POMDP, because historical observations must be included to successfully infer the next action. Formally, for a web agent task, if we represent the observation (including the screenshot and metadata of that webpage) at step $t$ as $o_{t}$, recent work represents the true state $s_{t}$ with $s_{t}=o_{t-n:t}$ (concatenation of the observations of last $n$ steps) (Bai et al., 2025b; Shen et al., 2025). This is called a windowed history, designed to fit into the context limit of the vision-language models. Formally, this modeling assumes the approximation
$\Pr\bigl(o_{t+1}|o_{\mathbf{1}:t},a_{\mathbf{1}:t-1},c\bigr)\approx\Pr\bigl(o_{t+1}|o_{\mathbf{t-n}:t},a_{\mathbf{t-n}:t-1},c\bigr),$ (2)
where $o$ is the vector that represents observation, and $c$ is the vector that represents task (Zhou et al., 2024; Shen et al., 2025; Bai et al., 2025b; QwenLM-Team, 2025). In words, these works treat observations of the last several steps as sufficient statistics of history for predicting future observations. In practice, the window size $n$ is usually set to $3$. For long-horizon tasks, this modeling over-simplifies the problem because the policy will frequently return to the same state. If the policy is not aware of its full history, it can produce repetitive sequences of actions because it is not aware of what it already finished earlier.
Previously, there have been clever designs in token-level RL (modeling a token as a step), e.g. RL from human feedback (RLHF), that mitigate this problem by directly represent $s_{t}$ with the representation of the token from the auto-regressive vision-language model, which automatically encodes information from all previous steps, naturally enforcing $s_{t}=o_{1:t}$. Is there a way we can bring this into sequence-level RL (modeling a sequence as a step)?
Practically, the problem with inputting full history observations $o_{1:t}$ to the VLM policy is that the context window of even the most advanced VLMs is not large enough, partially because image is too dense a representation for web agents tasks. Recent work like GLM-4.1v (Hong et al., 2025) addresses this limitation by introducing a memory as part of the tokens that the VLM should output in each step. Formally, the memory at step $t$, denoted by $m_{t}$, is a result of the observation without memory $\tilde{o}_{t}$, the memory from the previous step $m_{t-1}$, and the task instruction $c$, denoted as $m_{t}=\text{VLM}_{\theta}(\tilde{o}_{t},m_{t-1},c)$, and the observation at step $t$ can now be denoted as $o_{t}=(m_{t},\tilde{o}_{t},c)$. This design leads to a continuous compression of the memory based on the progress of the task, which resembles RLHF and Long-Short-Term Memory (LSTM). We practice this idea for our agent context management, as shown in Section E.1.
Appendix D Qualitative Examples
D.1 Example Rubrics
We show some example rubrics generated in Table 5. We can observe a general correlation between the instruction length of the task and the number of criteria in its rubrics.
Table 5: Examples of WebGym Tasks with Evaluation Rubrics
Task Evaluation Rubric
Find out the date of the next Baden Marathon event. Group 1: Find date of next Baden Marathon event
Facts:
- date of the next Baden Marathon event
Find the recommended daily feeding amount for an adult British Shorthair cat (over 12 months) on royalcanin.pl, assuming a target weight of 5kg and a moderate activity level. Group 1: Identify the target cat specifications
Facts:
- adult British Shorthair cat
- age over 12 months
- target weight of 5kg
- moderate activity level
Group 2: Find recommended daily feeding amount
Facts:
- recommended daily feeding amount for the specified cat on royalcanin.pl
Find the highest-rated mover in Honolulu to shift a vehicle and large appliances and who has virtual discussion options available in Yelp. Group 1: Identify highest-rated mover in Honolulu
Facts:
- mover is located in Honolulu
- mover is highest-rated
Group 2: Services provided by the mover
Facts:
- mover offers vehicle shifting services
- mover offers large appliances shifting services
Group 3: Communication options available
Facts:
- mover has virtual discussion options available
Group 4: Source of information
Facts:
- information is found on Yelp
D.2 Comparing Rubric-based Evaluation with Task-based Evaluation
We find many cases where rubric-based evaluation is able to produce significantly more comprehensive evaluation that checks details of each criteria of the task, which leads to many wins over using only the task description. One example is shown in Figure 12. The human evaluation for this example is NOT SUCCESS.
Figure 12: Trajectory for the comparison between vanilla evaluation and rubric-based evaluation.
Task description-based evaluation.
  • Verify criterion a: The response mentions the food groups eliminated during the Original Whole30 program: added sugar, alcohol, grains, legumes, dairy, and baked goods, pasta, cereal, chips, and fries. The screenshots provide evidence of these eliminations as listed on the Whole30 website. Additionally, the reasons mentioned in the response for eliminating these groups, such as identifying food sensitivities and improving relationship with food, are also corroborated by the screenshots, particularly those discussing the philosophy of Whole30.
  • Verify criterion b: The task asked for the food groups strictly eliminated during the Original Whole30 program and the reasons for their elimination according to the program’s philosophy. The agent’s response matches the task requirements accurately, listing the correct food groups and providing the reasons as asked, based on the information available in the screenshots.
  • Verdict: SUCCESS
Rubric-based evaluation.
  • There are 2 groups for this rubric in total.
  • Group 1: Identify food groups eliminated during the Original Whole30 program
    • Description: Identify food groups eliminated during the Original Whole30 program
    • Facts: Identify the food groups strictly eliminated during the Original Whole30 program
    • Verify criterion a: The response claims that the food groups strictly eliminated during the Original Whole30 program include added sugar, alcohol, grains, legumes, dairy, and items such as baked goods, pasta, cereal, chips, and fries. The screenshots clearly show a list of eliminated food groups that includes added sugar (real or artificial) alongside the other listed categories. Therefore, the response’s identification of eliminated food groups is supported by the screenshots.
    • Verify criterion b: This group requires identifying the food groups strictly eliminated during the Original Whole30 program. The response provides a concrete list of eliminated groups (including added sugar), which directly satisfies the identification requirement.
    • Verdict: SUCCESS
  • Group 2: Reasons for eliminating each food group
    • Description: Reasons for eliminating each food group
    • Facts: Reasons for eliminating each food group according to the Original Whole30 program’s philosophy
    • Verify criterion a: While the screenshots list the eliminated food groups, they do not explicitly provide the program’s philosophy-based reasons for eliminating each group. The response offers general motivations (e.g., identifying sensitivities or resetting habits), but these statements are not directly verifiable from the screenshots as provided.
    • Verify criterion b: This group specifically asks for reasons for eliminating each food group according to the Original Whole30 philosophy. The response does not provide distinct, food-group-specific reasons (it stays general rather than explaining each eliminated group), and the reasons it does provide are not grounded in the screenshots. Therefore, the response does not meet the requirement of explaining why each food group is eliminated.
    • Verdict: NOT SUCCESS
  • Final verdict: NOT SUCCESS
Appendix E Prompts
E.1 Agent Action Prompt
The prompt we use to rollout trajectories is the agent action prompts shown below. Both the official prompt as provided in QwenLM-Team (2025) and the memory prompt that we use across §5 are included below.
Prompt: Agent Action Generation
E.2 Evaluator Prompt: Keypoint Detection
Prompt: Evaluator/Evaluation Criteria Generation
E.3 Evaluator Prompt: Per-criteria Evaluation
The evaluation system uses a two-criterion approach, each checked separately:
  • Criterion A is checked per-fact (each rubric/fact from task decomposition is evaluated separately).
  • Criterion B is checked once per task (verifies the agent’s final answer against screenshots).
E.3.1 Criterion A: Fact Verification
Prompt: Evaluator/Criterion A - Fact Verification
E.3.2 Criterion B: Response Verification / Anti-Hallucination
Prompt: Evaluator/Criterion B - Anti-Hallucination Check
E.4 Evaluator Prompt: Blocking Detection
Additionaly, we introduce a blocking detection step, which is used to identify if failures were due to anti-bot measures vs. agent errors. Those websites that frequently blocks the agent will be excluded from being sampled.
Prompt: Evaluator/Blocking Detection
E.5 Task Set Generation Prompt: Evaluation Criteria Proposal
Prompt: Task/Evaluation Criteria Generation
E.6 Task Set Generation Prompt: Task Proposal from Fact Groups
Prompt: Task/Task Generation from Selected Fact Groups
Appendix F Hyper-parameters
Our system consists of four main components: environment management, data collection (rollout), model training, and the agent model. The hyperparameters for each component are carefully tuned to balance performance, scalability, and sample efficiency. The final hyperparameters can be found in Table 6.
Table 6: Hyperparameters for All Experiments. Bolded option suggests an experimented optimality.
Component Hyperparameter Value
Environment (2 workers) Simulation CPU Server size 80
HTTP pool size (screenshot/execute) 256
HTTP pool size (metadata/navigate) 256
HTTP pool size (allocate/release) 4
Max vLLM sessions 80
Rollout Train tasks rollout size 512, 1024
Task Sampling uniform/ratio
Max steps (easy,medium,hard) (15,30,45), (10,20,30)
Interaction mode coordinates-only, set-of-marks
Temperature 1.0
Top-p 0.99
Top-k 2
Training Learning rate 1e-5, 5e-6, 1e-6
Effective batch size 3*4*8*2
Times each sample is trained $\approx 2$
Max gradient norm 1.0
Weight decay 0, 0.01
Warmup steps 30, 100
LR scheduler type constant_with_warmup, cosine
Samples to train per iteration 1200, 1800
Recency bias power 1, 2
Replay buffer capacity last 4, last 8
Cutoff length 4096, 8192, 16384
Gradient checkpointing False
bf16 precision True
Model Model type Qwen3-VL-8B
Variant Instruct, Thinking
Max new tokens 1024, 2048, 3072